Pipeline sessions branch merge testing, found major new inputs handling error
AQUA report presentation to the JVLA pipeline coordination meeting
Email discussions on ALMA dbservice, archive connection, Cycle 6, etc issues
plotting cal tables in plotms: CAS-8270 under validation for minor bugfixes, CAS-10900 for printing summary to log (a la MSSummary) under development
Working on CAS-10458: Vi/VB2 vb->visCubeCorrected() segfaults if no column. Improving unit tests on VI/VB2 while at that.
CAS-10644 : skipping duplicates in DATA and POINTING. The POINTING fix has been validated. Still working on the DATA changes.
CAS-10939 - importasdm error, exit code 256, Not a bug, closed (no data of the requested type existed in that ASDM).
CAS-10386 - Add new shared ALMA code to support 2 new ASDM tables in asdm2MS. The code has been copied to my sandbox and I've verified that test_importasdm still passes. Working on some example ASDMs from the VLA (pulsar table) to use to check this. Possibly also an ALMA ASDM with the new telcal-related table in it. There's a hint in the e-mail comments from the VLA that getting pulsar data to fill correctly may require additional asdm2MS changes.
importasdm (asdm2MS) silently skips Main rows when the binary file pointed to in that row is missing. Jeff remembers this as being a feature, not a bug. It may be time to reopen discussion on what should happen in that case. I'm still verifying that this is what happens.
Dirk Petry (cannot attend this time due to travel)
Problem detected last week in Measures table repository: updates didn't arrive at the repository. (This was detected by the automatic test cron job here at ESO.)
Multiple causes, partially masking each other: (1) accidental commit of older version by Tak replaced newer version, (2) power cut as ASTRON prevented update for one week, (3) server ("catapult") setup problem at NRAO
Communicated with Darrell who performed various fixes in the repository and elsewhere.
The automatic update system at ESO has been reset.
As of 26 Nov, ca. 15:00 h UT, all is fine again (both updates via update-data and direct rsync give the desired result; automatic update of this morning went fine; latest version of IERSpredict is 26 Nov 2017, available from the repo)
spent most of the end of the week understanding the /home/casa mirroring issues at the AOC
Added unit tests for cube tclean for an MS with a decreasing frequency order (CAS-10940). Also updated Spectral Imaging section in the plone documentation describing output cube image channel ordering for default width.
Discussion on possible use of the JPL-Horizons ephemeris tables in data/ephemerides/JPL-Horizons for ALMA control with Ralph. They will need much higher time sampling.
Removed limb-darkened disk model from cl tools model option as the current implementation will not work for directly creating image from the componentlist (CAS-10903)